### v1

# Verifying the Resonance Schemes of Unstable Particles at Lepton Colliders

Abstract
We propose practical ways of differentiating the various (Breit-Wigner, theoretical, and energy-dependent) resonance schemes of unstable particles at lepton colliders. First, the energy-dependent scheme can be distinguished from the other two by fitting the $Z$ lineshape scan and forward-backward asymmetries at LEP and future lepton colliders with the $Z$ mass $m_{Z}$, decay width $\Gamma_{Z}$, and coupling strength as fitting parameters. Although the Breit-Wigner and theoretical schemes work equally well, the scheme conversion requires the decay width $\Gamma_{Z}$ to scale inversely with $m_{Z}$ rather than the usual linear dependence from theoretical calculation. These contradicting behaviors can be used to distinguish the Breit-Wigner and theoretical schemes by the precision $Z$ measurements with single parameter ($m_{Z}$) fit at future lepton colliders. For the $WW$ threshold scan, its combination with the precise Fermi constant provides another way of distinguishing the Breit-Wigner and theoretical schemes.
Introduction – The description of unstable particles is an intriguing part of quantum and field theories. A commonly adopted parametrization is the so-called Breit-Wigner resonance Breit:1936zzb ,
$\mathcal{D}_{\rm BW}\sim\frac{1}{p^{2}-m^{2}_{\rm BW}+im_{\rm BW}\Gamma_{\rm BW}},$ (1)
whose magnitude squared $1/[(p^{2}-m^{2}_{\rm BW})^{2}+m^{2}_{\rm BW}\Gamma^{2}_{\rm BW}]$ follows a resonance shape with pole at exactly the mass, $p^{2}=m^{2}_{\rm BW}$. The Breit-Wigner (BW) scheme give a very intuitive understanding of the pole mass $m_{\rm BW}$. Since the propagator can enter concrete processes in various ways, we refrain from replacing $p^{2}$ by the Mandelstam variable $s$ that is usually for the $s$-channel processes to make it general. A review of more details can be found in Giacosa:2021mbz .
A physical particle should have definite dispersion relation, $E=\sqrt{\bm{p}^{2}+m^{2}}$ where $(E,\bm{p})$ is the 4-momentum and $m$ is mass, according to Einstein’s special relativity. With a real energy $E$, the particle state evolves into the infinite future with a time-dependent evolution phase $e^{-iEt}$. So a real energy $E$ associated with time evolution intrinsically describes a stable particle. On the other hand, an unstable particle develops an imaginary part $-\frac{i}{2}\Gamma$ that is attached to the real energy, $E\rightarrow E-\frac{i}{2}\Gamma$. Then the particle wave function $\sim e^{-iEt-\frac{1}{2}\Gamma t}$ keeps decreasing with time and the corresponding decay width (lifetime) is $\Gamma$ ($1/\Gamma$).
It is much more intuitive to define the exact value of decay width in the rest frame. In other words, the physical pole should actually happen at $E=m-\frac{i}{2}\Gamma\equiv\mu$. Considering the fact that a relativistic dispersion relation should be Lorentz invariant, the propagator contains $p^{2}$ in the denominator Willenbrock:1991hu ; Willenbrock:2022smq ,
$\mathcal{D}_{\rm th}\sim\frac{1}{p^{2}-\mu^{2}}=\frac{1}{p^{2}-m^{2}_{\rm th}+\frac{1}{4}\Gamma^{2}_{\rm th}+im_{\rm th}\Gamma_{\rm th}}.$ (2)
The pole $E=\sqrt{\bm{p}^{2}+\mu^{2}}$ then happens at $\mu$ in the rest frame with vanishing $\bm{p}$. Since this scheme comes from theoretical arguments and has the benefit of being gauge independent Sirlin:1991fd , we call it the theoretical (th) scheme.
Comparing with the Breit-Wigner resonance in Eq. (1), the theoretical scheme has an extra term $\frac{1}{4}\Gamma^{2}$. For a narrow resonance, $\Gamma\ll m$, the Breit-Wigner scheme is a good enough approximation. However, the $Z$ mass can shift by 8.5 MeV which seems to be already larger than the current precision $\Delta m_{Z}=2.1$ MeV Workman:2022ynf .
Another commonly used scheme for the $Z$ pole scan is the energy-dependent scheme Berends:1987bg ; Bardin:1988xt ; Bardin:1989di ,
$\mathcal{D}_{\rm ED}\sim\frac{1}{p^{2}-m^{2}_{\rm ED}+i\frac{p^{2}\Gamma_{\rm ED}}{m_{\rm ED}}},$ (3)
[…]
Since the propagator schemes would affect the interpretation of the experimental data almost everywhere, especially for the precision test of new physics beyond the Standard Model (SM) of particle physics, it is of great importance to testify them with real data which is still an open question. We would thoroughly explore the phenomenological consequences with the $Z$ lineshape and $WW$ threshold scans at lepton colliders. The precision measurement of Fermi constant $G_{F}$ can also provide an input for the $W$ mass. For convenience, we take the commonly-used Breit-Wigner scheme as anchor in our following discussions.
Breit-Wigner vs Energy-Dependent Schemes – The comparison between the Breit-Wigner scheme in Eq. (1) and its energy-dependent counterpart in Eq. (3) should be done at the amplitude squared level. This is because the energy-dependent scheme leads to,
$|\mathcal{D}_{\rm ED}|^{2}\sim\frac{\mathcal{Z}}{(p^{2}-\mathcal{Z}m^{2}_{\rm ED})^{2}+\mathcal{Z}^{2}m^{2}_{\rm ED}\Gamma^{2}_{\rm ED}},$ (4)
where $\mathcal{Z}\equiv 1/(1+\Gamma^{2}_{\rm ED}/m^{2}_{\rm ED})$ with the $p^{2}$ correction from the decay width term. Comparing with the traditional Breit-Wigner scheme $|\mathcal{D}_{\rm BW}|^{2}\sim 1/[(p^{2}-m^{2}_{\rm BW})^{2}+m^{2}_{\rm BW}\Gamma^{2}_{\rm BW}]$, the peak location and decay width shifts as,
$m_{\rm BW}=\sqrt{\mathcal{Z}}m_{\rm ED},\quad\mbox{and}\quad\Gamma_{\rm BW}=\sqrt{\mathcal{Z}}\Gamma_{\rm ED}.$ (5)
In addition, the overall $\mathcal{Z}$ factor can be absorbed into the field redefinition. Since each propagator connects two vertices or two fields, each side should receive a scaling factor of $\sqrt{\mathcal{Z}}$. Take the fermion gauge coupling vertex $V\bar{f}f$ with $V=Z$ or $W$ as example, the coupling strength $g$ should be rescaled/renormalized as $g_{\rm BW}=\mathcal{Z}^{1/4}g_{\rm ED}$.
The Breit-Wigner and energy-dependent schemes agree quite well with each other for the narrow width approximation ($\Gamma\ll m$). Taking the weak gauge bosons for illustration, the width is just 2.7% for both $Z$ ($m_{Z}=91.1876$ GeV vs $\Gamma_{Z}=2.4955$ GeV Workman:2022ynf ) and $W$ ($m_{W}=80.376$ GeV vs $\Gamma_{W}=2.195$ GeV Workman:2022ynf ). Consequently, the difference $\Gamma^{2}/2m^{2}$ between the two schemes is only at the level of $3.7\times 10^{-4}$ for $Z$ and $W$ corresponding to a shift of 34 MeV and 30 MeV in the $Z$ and $W$ masses, respectively. Although such correction seems quite tiny, the sensitivity at lepton colliders can reach the precision. For the $Z$ boson mass, the current precision that mainly comes from LEP is 2.1 MeV while it can further shrink to 0.1 MeV at CEPC CEPCPhysicsStudyGroup:2022uwl and FCC-ee Blondel:2021ema . It seems possible and is of great importance to use the in-situ $Z$ lineshape scan to test the resonant schemes for an unambiguous conclusion.
However, since the propagator amplitude squares $|\mathcal{D}_{\rm BW}|^{2}$ and $|\mathcal{D}_{\rm ED}|^{2}$ can be made equivalent, there is no essential difference between these two schemes. If only a single $Z$ propagator is involved, the resonance scan cannot tell the difference between the Breit-Wigner and energy-dependent schemes. Although the extracted $Z$ mass from the same data but with different schemes can differ by 34 MeV, differentiating them needs extra justification.
Fortunately, the $Z$ resonance at lepton colliders involves not just the $Z$ propagator but also the virtual photon ($\gamma$) mediation. Being massless, the photon is stable with vanishing decay width. In other words, the issue of resonance scheme appears only for the $Z$ propagator. But the two contributions $\mathcal{M}_{Z}$ and $\mathcal{M}_{\gamma}$ can interfere with each other to give an interference term, $\mathcal{M}_{Z}\mathcal{M}^{*}_{\gamma}$,
$(\mathcal{M}_{Z}\mathcal{M}^{*}_{\gamma})_{\rm BW}$ $\propto\frac{g^{2}_{Z,\rm BW}e^{2}(s-m^{2}_{\rm BW})}{\left[(s-m^{2}_{\rm BW})^{2}+m^{2}_{\rm BW}\Gamma^{2}_{\rm BW}\right]s},$ (6a)
$(\mathcal{M}_{Z}\mathcal{M}^{*}_{\gamma})_{\rm ED}$ $\propto\frac{\mathcal{Z}g^{2}_{Z,\rm ED}e^{2}(s-m^{2}_{\rm ED})}{\left[(s-\mathcal{Z}m^{2}_{\rm ED})^{2}+\mathcal{Z}^{2}m^{2}_{\rm ED}\Gamma^{2}_{\rm ED}\right]s},$ (6b)
where $g_{Z}$ and $e$ stands for the weak and electric couplings. Note that the $\mathcal{Z}$ factor only affects the weak coupling $g_{Z}$ while the electric charge $e$ attached with photon can be fixed by various electromagnetic processes and hence should remain the same when discussing the resonance schemes. Although the pole structure in the denominators of Eq. (6) can be adjusted to be the same according to Eq. (5), the linear term $s-m^{2}_{\rm BW}$ and $s-m^{2}_{\rm ED}$ in the numerators cannot match simultaneously. In other words, it is not possible to match both the squared $|\mathcal{M}_{Z}|^{2}$ and interference $\mathcal{M}_{Z}\mathcal{M}^{*}_{\gamma}$ terms at the same time. The difference between the Breit-Wigner and energy-dependent schemes can arise off the $Z$ pole with nonvanishing linear terms.
For the total cross section with unpolarized $e^{+}e^{-}$ beams, both left- and right-handed electrons can contribute. The $g^{2}_{Z}$ term for the interference term then takes the form as Workman:2022ynf ,
$\sigma_{Z\gamma}=\frac{\alpha Q_{e}Q_{f}}{6}\frac{(g_{eL}+g_{eR})(g_{fL}+g_{fR})(s-m_{Z}^{2})}{(s-m_{Z}^{2})^{2}+\Gamma_{Z}^{2}m_{Z}^{2}},$ (7)
for the Breit-Wigner scheme and similarly for other schemes. The interference term is modulated by $g_{eL}+g_{eR}$ for the initial $e^{+}e^{-}$ and $g_{fL}+g_{fR}$ for the final $f\bar{f}$ fermions. Note that the charged leptons have $g_{\ell L}=-\frac{1}{2}+s^{2}_{w}$ and $g_{\ell R}=s^{2}_{w}$ with their prefactor $e/(c_{w}s_{w})$, where $s_{w}\equiv\sin\theta_{w}$ ($c_{w}\equiv\cos\theta_{w}$) is the sine (cosine) function of the weak mixing angle $\theta_{w}$. With $s^{2}_{w}\approx 0.223$, the combination $g_{\ell L}+g_{\ell R}=-\frac{1}{2}+2s^{2}_{w}$ is quite suppressed. In addition, the interference term receives another suppression from $s-m^{2}_{Z}$. The interference term is intrinsically subdominant in the total cross section of unpolarized $e^{+}e^{-}$ collision.
To make the interference term more explicit, it is desirable to have either polarized beams for the initial $e^{+}e^{-}$ or forward-backward asymmetry for the final-state fermions. The beam polarization would replace the electron couplings with $f_{L}g_{eL}+f_{R}g_{eR}$ where $f_{L}$ and $f_{R}$ are the beam polarization fractions while the final-state factor $g_{fL}+g_{fR}$ still remains suppressed for the leptonic final states. For comparison, the forward-background asymmetry can fully break the suppression since the cross section difference Workman:2022ynf ; ALEPH:2006bhb ,
$\sigma_{F}-\sigma_{B}$ $\propto\frac{3s}{64\pi}(g^{2}_{eL}-g^{2}_{eR})(g^{2}_{fL}-g^{2}_{fR})$ (8)
$+\frac{3Q_{e}Q_{f}\alpha(s-M^{2}_{Z})}{8}\left(g_{eL}-g_{eR}\right)\left(g_{fL}-g_{fR}\right),$
is intrinsically modulated by $g_{eL}-g_{eR}$ and $g_{fL}-g_{fR}$ both having a magnitude of $1/2$. Comparing with the interference term in the second line, the $|\mathcal{M}_{Z}|^{2}$ term in the first line receives at least $g_{eL}+g_{eR}$ suppression instead.
As mentioned earlier, it is impossible to match the Breit-Wigner and energy-dependent schemes in the presence of both the $Z$ contribution $|\mathcal{M}_{Z}|^{2}$ and the $Z$-$\gamma$ mediation. So we take into consideration both the hadronic cross section $\sigma_{\rm had}$ that mainly provides the $|\mathcal{M}_{Z}|^{2}$ contribution and the forward-backward asymmetry $A^{f}_{\rm FB}$ for the interference term,
$\sigma_{\rm had}\equiv\sum_{q}\sigma^{q}_{F}+\sigma^{q}_{B},\quad A^{f}_{FB}\equiv\frac{\sigma^{f}_{F}-\sigma^{f}_{B}}{\sigma^{f}_{F}+\sigma^{f}_{B}}.$ (9)
The previous precision measurements at LEP can already reach the $10^{-3}$ level Workman:2022ynf ; RothBook . One may fit those $Z$ lineshape data from the four detectors ALEPH ALEPH:1999smx ; ALEPH:2001pzx , DELPHI DELPHI:2000wje ; DELPHI:2003fml , L3 L3:2000vgx ; L3:1998bss , and OPAL OPAL:2000ufp ; OPAL:2003pfe at LEP with both the $Z$ mass $m_{Z}$ and decay width $\Gamma_{Z}$ as independent parameters. In addition, we also treat the weak mixing angle $s^{2}_{w}$ as an extra fitting parameters. This fitting can be done for either scheme to produce its own $\chi^{2}_{\rm min}$
$\Delta\chi^{2}_{\rm min}$ ALEPH DELPHI L3 OPAL Combined
$\sigma_{\rm had}$ $0.106$ $0.0083$ $0.0127$ $0.0529$ $0.183$
$\sigma_{\rm had}+A^{\mu}_{FB}$ 8.34 10.1 5.70 14.1 37.9
$\sigma_{\rm had}+A^{b,c}_{FB}$ 1.19 0.878 0.010 0.734 2.97
$\sigma_{\rm had}+A^{\mu,b,c}_{FB}$ 8.42 9.67 5.58 13.3 37.0
Table 1: The sensitivity of distinguishing the Breit-Wigner and energy-dependent schemes with the existing LEP data. The forward-backward asymmetry of a charm quark is not considered at L3 due to its lack of statistics ().
and the difference $\Delta\chi^{2}_{\rm min}\equiv|\chi^{2}_{\rm ED,\,min}-\chi^{2}_{\rm BW,\,min}|$ as summarized in Tab. 1 characterizes the capability of LEP in distinguishing the two schemes. Although the $Z$ lineshape data at various energy points have already been included, the fit with $\sigma_{\rm had}$ alone does not have much sensitivity as expected. Only when the $A^{f}_{FB}$ measurements ($f=\mu,b,c$) are also included, the combined sensitivity can reach $\Delta\chi^{2}_{\rm min}=37$.
At CEPC, a $100\,{\rm ab}^{-1}$ of integrated luminosity is projected at the $Z$ pole ($\sqrt{s}\simeq 91.2\,{\rm GeV}$) to generate $3\times 10^{12}$ events CEPCStudyGroup:2018ghi ; CEPCPhysicsStudyGroup:2022uwl . The dominant uncertainty at CEPC comes from the integrated luminosity, $\delta\mathcal{L}/\mathcal{L}=5\times 10^{-5}$ CEPCPhysicsStudyGroup:2022uwl while its statistical uncertainty $1/\sqrt{N}$ can reach the $10^{-7}$ level. Those off-$Z$ pole points (87.9, 90.2, 92.2, 94.3) GeV would have $1\,{\rm ab}^{-1}$ each. Since the forward-backward asymmetry $A^{f}_{FB}$ in Eq. (9) is defined as ratio of events $N_{F/B}\propto\mathcal{L}\times\sigma_{F/B}$, the luminosity $\mathcal{L}$ and its uncertainty factorize out. Then the statistical uncertainty, $(\delta A_{FB})^{2}\simeq 4N_{F}N_{B}/(N_{F}+N_{B})^{3}$, dominates.
Our fit $\chi^{2}\equiv\sum_{i}(\mathcal{O}^{i}_{\rm BW}-\mathcal{O}^{i}_{\rm ED})^{2}/(\delta\mathcal{O}^{i})^{2}$ is done for the hadronic cross section $\sigma_{\rm had}$, the forward-backward asymmetries $A^{f}_{\rm FB}$, and their combination. The observable pseudo-data $\mathcal{O}^{i}_{\rm ED}$ and the corresponding uncertainty $\delta\mathcal{O}^{i}$ are estimated in the energy-dependent scheme with the PDG values Workman:2022ynf of $m_{Z}=91.1876$ GeV, $\Gamma_{Z}=2.4955$ GeV, and $s^{2}_{w}=0.223046$. These pseudo-data are then fit by the Breit-Wigner scheme predictions of $\mathcal{O}^{i}_{\rm BW}$ with 3 fitting parameters $m_{Z}$, $\Gamma_{Z}$, and $s^{2}_{w}$. If either $\sigma_{\rm had}$ or the forward-backward asymmetries $A^{q}_{FB}$ ($q=u,d,s,c,b$) are included in the analysis, there is almost no sensitivity at all. However, their combination can significantly boost to $\chi^{2}_{\rm min}(\sigma_{\rm had}+A^{q}_{\rm FB})=490$. Further including $A^{\mu}_{\rm FB}$, $\chi^{2}_{\rm min}$ jumps to $5.1\times 10^{5}$. Comparing with the LEP result, the $\Delta\chi^{2}_{\rm min}$ increase by a factor of $1.3\times 10^{4}$, being smaller than the $3\times 10^{5}$ times increase in event rates, is reasonable.
Breit-Wigner vs Theoretical Schemes – Between the Breit-Wigner scheme in Eq. (1) and the theoretical one in Eq. (2), the only difference is that the pole position shifts, $m^{2}_{\rm BW}\rightarrow m^{2}_{\rm th}-\frac{1}{4}\Gamma^{2}_{\rm th}$, while the imaginary part remains the same, $m_{\rm BW}\Gamma_{\rm BW}=m_{\rm th}\Gamma_{\rm th}$. Consequently, the two sets of parameters scale as,
$\frac{m_{\rm th}}{m_{\rm BW}}=\frac{\Gamma_{\rm BW}}{\Gamma_{\rm th}}\simeq 1+\frac{\Gamma^{2}_{Z}}{8m^{2}_{Z}}\equiv R_{Z}.$ (10)
The relative correction $9.2\times 10^{-5}$ is smaller than the difference between the Breit-Wigner and energy-dependent schemes by a factor of 4. Since the ratio $\Gamma^{2}_{Z}/m^{2}_{Z}$ is already a small number, there is no need to distinguish between the values of $\Gamma_{Z}$ and $m_{Z}$ in the Breit-Wigner or theoretical schemes. So we omit the scheme labels for such ratios hereafter.
The equivalence between these two schemes can be achieved at the amplitude level already, in contrast to the situation between the Breit-Wigner and energy-dependent schemes that the exact match can only happen at the amplitude squared level $|\mathcal{M}_{Z}|^{2}$. In other words, the Breit-Wigner and theoretical schemes share higher similarity. If the $m_{Z}$ and $\Gamma_{Z}$ are used as two free parameters, the Breit-Wigner and theoretical schemes can fit the $Z$ lineshape equally well Willenbrock:2022smq . Due to the same similarity, the $Z$ lineshape scan and forward-backward asymmetries discussed in the previous section can also apply to tell the difference between the theoretical and energy-dependent schemes.
The only way for distinguishing the Breit-Wigner and theoretical schemes probably lies in making theoretical connection between $m_{Z}$ and $\Gamma_{Z}$. As shown in Eq. (10), the mass and decay width are inversely proportional to each other. When extracting the decay width from data, the decay width takes a smaller value for a heavier $Z$ boson, $\Gamma_{Z}\propto 1/m_{Z}$, using either the Breit-Wigner or the theoretical scheme. However, the naive expectation with dimensional analysis tells us that the decay width of an unstable particle is typically proportional to its mass, $\Gamma_{Z}\propto m_{Z}$ for the $Z\rightarrow f\bar{f}$ decay,
$\Gamma_{Z}=\frac{m_{Z}}{24\pi}\beta_{f}\left[g^{2}_{L}+g^{2}_{R}-(g^{2}_{L}-6g_{L}g_{R}+g^{2}_{R})\frac{m^{2}_{f}}{m^{2}_{Z}}\right],$ (11)
where $m_{f}$ is the final-state fermion $f$ mass while $g_{L}$ and $g_{R}$ are the left- and right-handed couplings. To keep formula compact, we have defined the phase space factor $\beta_{f}\equiv\sqrt{1-4m^{2}_{f}/m^{2}_{Z}}$. The full decay width $\Gamma_{Z}$ is a total contribution from quarks, charged leptons, and neutrinos.
This mismatch between the extracted $\Gamma_{Z}$ from data and its theoretically predicted value can be experimentally tested. To be more specific, once the $Z$ lineshape fit is done with the Breit-Wigner scheme for example, the best-fit value of $m_{\rm BW}$ can be used to predict $\Gamma_{Z}$ as Eq. (11). Since there is no real data yet, the central values are taken as the current $m_{Z}=91.1876$ GeV Workman:2022ynf and the derived $\Gamma_{Z}=2.47042$ GeV with Eq. (11). For such setup, the Breit-Wigner scheme with one fitting parameter $m_{\rm BW}$ can perfectly fit the pseudo-data. For the theoretical scheme, we take $m_{\rm th}$ as the fitting parameter and $\Gamma_{\rm th}(m_{\rm th})$ is consequently predicted as function of $m_{\rm th}$ according to Eq. (11). However the fitting should still be be done after conversion to the Breit-Wigner scheme,
$\chi^{2}_{Z}\equiv\left(\frac{m_{\rm th}/R_{Z}-m_{Z}}{\Delta m_{Z}}\right)^{2}+\left(\frac{\Gamma_{\rm th}R_{Z}-\Gamma_{Z}}{\Delta\Gamma_{Z}}\right)^{2},$ (12)
using the scheme conversion factor $R_{Z}$ in Eq. (10).
Comparing with the $9.2\times 10^{-5}$ correction of Eq. (10), a thumb rule is that both precisions $\Delta m_{Z}\lesssim 8.4$ MeV and $\Delta\Gamma_{Z}\lesssim 0.023$ MeV are necessary to have sizable sensitivity. The current sensitivities $\Delta m_{Z}=2.1$ MeV and $\Delta\Gamma_{Z}=2.3$ MeV Workman:2022ynf are not good enough. However, with the projected sensitivity $\Delta m_{Z}=0.1$ MeV and $\Delta\Gamma_{Z}=0.025$ MeV at CEPC CEPCPhysicsStudyGroup:2022uwl and FCC-ee Blondel:2021ema , $\chi^{2}_{Z,\rm min}$ can reach 326. For the ILC-GigaZ projection of $\Delta m_{Z}=0.2$ MeV and $\Delta\Gamma_{Z}=0.12$ MeV Belloni:2022due , the projected sensitivity can also touch $\chi^{2}_{Z,\rm min}=14.3$.
Fermi Constant and $WW$ Threshold Scan – The Fermi constant $G_{F}$ has been measured very precisely in processes such as muon decay Workman:2022ynf . With the Breit-Wigner scheme, the Fermi constant usually takes the form as,
$G_{F,\rm BW}\approx\frac{g^{2}}{4\sqrt{2}m^{2}_{W,\rm BW}}\left(1-\frac{\overline{\Gamma}^{2}_{W}}{2m^{2}_{W}}\right),$ (13)
where $\overline{\Gamma}_{W}$ is not necessarily the full $W$ decay width which will be further elaborated below. For comparison, the theoretical scheme in Eq. (2) would lead to,
$G_{F,\rm th}\approx\frac{g^{2}}{4\sqrt{2}m^{2}_{W,\rm th}}\left(1-\frac{\overline{\Gamma}^{2}_{W}}{4m^{2}_{W}}\right).$ (14)
With the same measurement of the Fermi constant, the scheme conversion resembles Eq. (10) but with the $W$ mass $m_{W}$ and reduced decay with $\overline{\Gamma}_{W}$.
However, the $W$ virtuality $p^{2}$ in the muon decay process is smaller than the muon mass squared $m^{2}_{\mu}$, $p^{2}<m^{2}_{\mu}$. The reduced $W$ decay width $\overline{\Gamma}_{W}\approx\frac{2}{5}\Gamma_{W}$ should only consider the partial contributions Giacosa:2021mbz from $W\rightarrow e\nu_{e}$ and $ud$. In addition, the energy dependence of the decay with $\overline{\Gamma}_{W}(p^{2})=(p^{2}/m^{2}_{W})\overline{\Gamma}_{W}$ Berends:1987bg ; Bardin:1988xt ; Bardin:1989di gives further suppression, $p^{2}/m^{2}_{W}\leq m^{2}_{\mu}/m^{2}_{W}$. Altogether, the reduced decay width $\overline{\Gamma}(p^{2})\lesssim 7\times 10^{-7}\Gamma_{W}$ becomes negligibly small. In other words, the Fermi constant, $G_{F}\approx g^{2}/4\sqrt{2}m^{2}_{W}$, would force the $W$ mass to the same value regardless of the scheme.
However, a single data input from the Fermi constant is not enough to distinguish the two interpretations with the Breit-Wigner and theoretical schemes. It is necessary to find an independent measurement of the $W$ mass $m_{W}$ and decay width $\Gamma_{W}$,
$\Gamma_{W}=\frac{m_{W}}{24\pi}g^{2}\beta_{12}\left[1-\frac{m^{2}_{1}+m^{2}_{2}}{2m^{2}_{W}}-\frac{(m^{2}_{1}-m^{2}_{2})^{2}}{2m^{4}_{W}}\right],$ (15)
where the phase space factor is $\beta_{12}\equiv\beta_{+}\beta_{-}$ with $\beta_{\pm}\equiv\sqrt{1-(m_{1}\pm m_{2})^{2}/m^{2}_{W}}$ and $m_{1,2}$ are the final-state fermion masses. Considering the measured sizes, $m_{W}=80.377$ GeV and $\Gamma_{W}=2.085$ GeV Workman:2022ynf , the correction $8.1\times 10^{-5}$ is larger than the Fermi constant precision Workman:2022ynf .
At the future lepton colliders, the $WW$ threshold scan can provide precise measurement of $m_{W}$ and $\Gamma_{W}$. For example, the projected precision can reach $\Delta m_{W}=0.5$ MeV and $\Delta\Gamma_{W}=2.0$ MeV at CEPC CEPCPhysicsStudyGroup:2022uwl , $\Delta m_{W}=0.4$ MeV and $\Delta\Gamma_{W}=1.2$ MeV at FCC Blondel:2021ema , as well as $\Delta m_{W}=2.4$ MeV and $\Delta\Gamma_{W}=2$ MeV at ILC deBlas:2022ofj . Following the same procedure as the $Z$ mass and decay width fit in the previous section, a single-parameter fit can also be done for the $WW$ threshold scan that provides a simultaneous determination of $m_{W}$ and $\Gamma_{W}$, in addition to the Fermi constant,
$\chi^{2}_{W}$ $\equiv\left(\frac{m_{W,\rm th}/R_{W}-m_{W}}{\Delta m_{W}}\right)^{2}+\left(\frac{\Gamma_{W,\rm th}R_{W}-\Gamma_{W}}{\Delta\Gamma_{W}}\right)^{2}$
$+\left[\frac{G_{F,\rm th}-G_{F}}{\Delta G_{F}}\right]^{2},$ (16)
where $R_{W}$ takes the similar form as $R_{Z}$ in Eq. (10). For the central values we take the current $m_{W}=80.376$ GeV Workman:2022ynf and the derived $\Gamma_{W}=2.04501$ from Eq. (15), in addition to $G_{F}=1.166\,378\,8\times 10^{-5}\,\mbox{GeV}^{-2}$ Workman:2022ynf . The weak coupling is then fixed to $g=0.652986$ by Eq. (13) in the Breit-Wigner scheme. The sensitivity can reach $\Delta\chi^{2}_{W,\rm min}=169$, 263, and 7.3 at CEPC, FCC-ee, and ILC, respectively. By fitting the $WW$ threshold scan data, the extract $W$ mass values from the Breit-Wigner and theoretical schemes should also follow the scheme conversion in Eq. (10). However, the precisely measurement Fermi constant tends to give exactly the same value for $m_{W}$ in both schemes. This contradiction gives the distinguishing power.
Conclusions – Although various resonance schemes of unstable particles have been extensively studied, using electroweak precision measurement data to verify which scheme is correct is still an open question. With simple phenomenological $\chi^{2}$ fit, we found that the combination of $Z$ lineshape scan and the forward-backward asymmetry measurements at the $Z$ pole can tell the difference between the energy-dependent scheme and its Breit-Wigner (theoretical) counterpart with the existing data from LEP. Even higher sensitivity can be expected at the future CEPC and FCC-ee. For the remaining Breit-Wigner and theoretical schemes, they were thought to be fully equivalent and give exactly the same resonance shape, we also found a way of using the opposite scaling behaviors of the extracted decay width from data, $\Gamma_{Z}\propto 1/m_{Z}$, and the theoretically predicted one, $\Gamma_{Z}\propto m_{Z}$, to set a realistic approach for future lepton colliders. For the $WW$ threshold scan, the supplementation from the Fermi constant is necessary. More detailed simulation and data analysis from the experimental side is necessary. In summary, the resonance shape alone cannot distinguish its various schemes but needs to involve the forward-backward asymmetries and the scaling behaviors of decay width. These provide realistic tests of an important aspect of quantum and field theories.
Acknowledgements
The authors are grateful to Manqi Ruan and Junping Tian for providing useful information. SFG and UM are supported by the National Natural Science Foundation of China (12425506, 12375101, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604). ZQ is supported by the National Natural Science Foundation of China (1240050404) and the HZNU start-up fund. SFG is also an affiliate member of Kavli IPMU, University of Tokyo.

### v2

@@
 Introduction – The description of unstable particles is an intriguing part of quantum and field theories. A commonly adopted parametrization is the so-called Breit-Wigner resonance Breit:1936zzb ,
-$\mathcal{D}_{\rm BW}\sim\frac{1}{p^{2}-m^{2}_{\rm BW}+im_{\rm BW}\Gamma_{\rm BW}},$ (1)
-whose magnitude squared $1/[(p^{2}-m^{2}_{\rm BW})^{2}+m^{2}_{\rm BW}\Gamma^{2}_{\rm BW}]$ follows a resonance shape with pole at exactly the mass, $p^{2}=m^{2}_{\rm BW}$.
+$\mathcal{D}_{\rm BW}\sim\frac{1}{p^{2}-m^{2}_{\rm BW}+im_{\rm BW}\Gamma_{\rm BW}}$ (1)
+in natural units ($\hbar=c=1$), whose magnitude squared $1/[(p^{2}-m^{2}_{\rm BW})^{2}+m^{2}_{\rm BW}\Gamma^{2}_{\rm BW}]$ follows a resonance shape with pole at exactly the mass, $p^{2}=m^{2}_{\rm BW}$.
 The Breit-Wigner (BW) scheme give a very intuitive understanding of the pole mass $m_{\rm BW}$.
 Since the propagator can enter concrete processes in various ways, we refrain from replacing $p^{2}$ by the Mandelstam variable $s$ that is usually for the $s$-channel processes to make it general.
-A review of more details can be found in Giacosa:2021mbz .
-A physical particle should have definite dispersion relation, $E=\sqrt{\bm{p}^{2}+m^{2}}$ where $(E,\bm{p})$ is the 4-momentum and $m$ is mass, according to Einstein’s special relativity.
+A review of more details can be found in Bohm:2004zi ; Giacosa:2021mbz .
+A physical particle should have definite dispersion relation, $E=\sqrt{\bm{p}^{2}+m^{2}}$ where $(E,\bm{p})$ is the 4-momentum and $m$ is mass.
 With a real energy $E$, the particle state evolves into the infinite future with a time-dependent evolution phase $e^{-iEt}$.
@@
 It is much more intuitive to define the exact value of decay width in the rest frame.
-In other words, the physical pole should actually happen at $E=m-\frac{i}{2}\Gamma\equiv\mu$.
+In other words, the physical pole should actually happen at $E=m-\frac{i}{2}\Gamma\equiv\mu$, which is consistent with the linear approximation for the non-relativistic limit Dalitz:1961dx ; Dalitz:1963xk .
 Considering the fact that a relativistic dispersion relation should be Lorentz invariant, the propagator contains $p^{2}$ in the denominator Willenbrock:1991hu ; Willenbrock:2022smq ,
@@
 The pole $E=\sqrt{\bm{p}^{2}+\mu^{2}}$ then happens at $\mu$ in the rest frame with vanishing $\bm{p}$.
-Since this scheme comes from theoretical arguments and has the benefit of being gauge independent Sirlin:1991fd , we call it the theoretical (th) scheme.
+Since this scheme comes from theoretical arguments and has the benefit of being gauge independent Sirlin:1991fd ; Papavassiliou:1995fq , we call it the theoretical (th) scheme.
 Comparing with the Breit-Wigner resonance in Eq.
@@
 For a narrow resonance, $\Gamma\ll m$, the Breit-Wigner scheme is a good enough approximation.
-However, the $Z$ mass can shift by 8.5 MeV which seems to be already larger than the current precision $\Delta m_{Z}=2.1$ MeV Workman:2022ynf .
+By solving the equations, $m^{2}_{\rm BW}=m^{2}_{\rm th}-\frac{1}{4}\Gamma^{2}_{\rm th}$ and $m_{\rm BW}\Gamma_{\rm BW}=m_{\rm th}\Gamma_{\rm th}$, however, the $Z$ mass can shift by $m_{\rm th}-m_{\rm BW}=8.5$ MeV which seems to be already larger than the current precision $\Delta m_{Z}=2.1$ MeV Workman:2022ynf .
 Another commonly used scheme for the $Z$ pole scan is the energy-dependent scheme Berends:1987bg ; Bardin:1988xt ; Bardin:1989di ,
@@
 The Breit-Wigner and energy-dependent schemes agree quite well with each other for the narrow width approximation ($\Gamma\ll m$).
-Taking the weak gauge bosons for illustration, the width is just 2.7% for both $Z$ ($m_{Z}=91.1876$ GeV vs $\Gamma_{Z}=2.4955$ GeV Workman:2022ynf ) and $W$ ($m_{W}=80.376$ GeV vs $\Gamma_{W}=2.195$ GeV Workman:2022ynf ).
-Consequently, the difference $\Gamma^{2}/2m^{2}$ between the two schemes is only at the level of $3.7\times 10^{-4}$ for $Z$ and $W$ corresponding to a shift of 34 MeV and 30 MeV in the $Z$ and $W$ masses, respectively.
+Taking the weak gauge bosons for illustration, the width is just 2.7% of the mass for both $Z$ ($m_{Z}=91.1876\pm 0.0021$ GeV vs $\Gamma_{Z}=2.4955\pm 0.0023$ GeV at LEP 1 Workman:2022ynf ; ALEPH:2005ab ) and $W$ ($m_{W}=80.376\pm 0.033$ GeV vs $\Gamma_{W}=2.195\pm 0.83$ GeV at LEP 2 Workman:2022ynf ; ALEPH:2013dgf ).
+Consequently, the relative difference $\Gamma^{2}/2m^{2}$ between the two schemes is only at the level of $3.7\times 10^{-4}$ for the $Z$ and $W$ bosons.
+This corresponds to shifts of 34 MeV and 30 MeV in the $Z$ and $W$ masses, respectively, and correspondingly a shift of 0.93 MeV and 0.82 MeV in their decay widths.
 Although such correction seems quite tiny, the sensitivity at lepton colliders can reach the precision.
+[^1] The recent measurements of $m_{W}$ and $\Gamma_{W}$ at the LHC achieved a precision of $\Delta m_{W}=15.9$ MeV and $\Delta\Gamma_{W}=47$ MeV from the ATLAS Collaboration ATLAS:2024erm .
+Similarly, the CMS Collaboration reached $\Delta m_{W}=9.9$ MeV CMS:2024nau while $\Delta\Gamma_{W}$ was not explicitly reported.
+To distinguish between the two resonance schemes, high precision in both the mass and decay width measurements is essential.
+Despite having achieved sufficient precision in mass measurements, it is still challenging for hadron colliders to differentiate between the resonance schemes.
 For the $Z$ boson mass, the current precision that mainly comes from LEP is 2.1 MeV while it can further shrink to 0.1 MeV at CEPC CEPCPhysicsStudyGroup:2022uwl and FCC-ee Blondel:2021ema .
@@
 Comparing with the LEP result, the $\Delta\chi^{2}_{\rm min}$ increase by a factor of $1.3\times 10^{4}$, being smaller than the $3\times 10^{5}$ times increase in event rates, is reasonable.
+Notice that the soft real and radiative virtual corrections can have significant effect on the forward-backward asymmetry $A^{f}_{\rm FB}$ Greco:1980mh .
+Especially, the value of $A^{f}_{\rm FB}$ reduces above the $Z$ pole which would make the relative difference between different resonance schemes even larger.
+As a conservative estimation of the sensitivity, we adopt the tree-level calculation and leave the more sophisticated soft real and radiative virtual corrections for future study.
 Breit-Wigner vs Theoretical Schemes – Between the Breit-Wigner scheme in Eq.
@@
 Due to the same similarity, the $Z$ lineshape scan and forward-backward asymmetries discussed in the previous section can also apply to tell the difference between the theoretical and energy-dependent schemes.
-The only way for distinguishing the Breit-Wigner and theoretical schemes probably lies in making theoretical connection between $m_{Z}$ and $\Gamma_{Z}$.
-As shown in Eq.
-(10), the mass and decay width are inversely proportional to each other.
-When extracting the decay width from data, the decay width takes a smaller value for a heavier $Z$ boson, $\Gamma_{Z}\propto 1/m_{Z}$, using either the Breit-Wigner or the theoretical scheme.
+The only way for distinguishing the Breit-Wigner and theoretical schemes probably lies in making extra connection between $m_{Z}$ and $\Gamma_{Z}$.
+Eq.
+(10) requires the decay width to be inversely proportional to the mass, $\Gamma_{Z}\propto 1/m_{Z}$.
+This relationship arises because the observable quantity as data is the imaginary part of the self-energy function, ${\rm Im}\,\Pi_{Z}=m_{Z}\Gamma_{Z}$.
+When extracting the decay width from data, a heavier $Z$ boson corresponds to a smaller decay width, $\Gamma_{Z}\propto 1/m_{Z}$, regardless whether the Breit-Wigner or the theoretical scheme is used.
 However, the naive expectation with dimensional analysis tells us that the decay width of an unstable particle is typically proportional to its mass, $\Gamma_{Z}\propto m_{Z}$ for the $Z\rightarrow f\bar{f}$ decay,
@@
 (11).
-However the fitting should still be be done after conversion to the Breit-Wigner scheme,
+However the fitting should still be done after conversion to the Breit-Wigner scheme,
 $\chi^{2}_{Z}\equiv\left(\frac{m_{\rm th}/R_{Z}-m_{Z}}{\Delta m_{Z}}\right)^{2}+\left(\frac{\Gamma_{\rm th}R_{Z}-\Gamma_{Z}}{\Delta\Gamma_{Z}}\right)^{2},$ (12)
@@
 Comparing with the $9.2\times 10^{-5}$ correction of Eq.
-(10), a thumb rule is that both precisions $\Delta m_{Z}\lesssim 8.4$ MeV and $\Delta\Gamma_{Z}\lesssim 0.023$ MeV are necessary to have sizable sensitivity.
+(10), a thumb rule is that both precisions $\Delta m_{Z}\lesssim 8.4$ MeV and $\Delta\Gamma_{Z}\lesssim 0.23$ MeV are necessary to have sizable sensitivity.
 The current sensitivities $\Delta m_{Z}=2.1$ MeV and $\Delta\Gamma_{Z}=2.3$ MeV Workman:2022ynf are not good enough.
@@
 For the ILC-GigaZ projection of $\Delta m_{Z}=0.2$ MeV and $\Delta\Gamma_{Z}=0.12$ MeV Belloni:2022due , the projected sensitivity can also touch $\chi^{2}_{Z,\rm min}=14.3$.
+We have also tried implementing the energy-dependent width for both the Breit-Wigner and theoretical schemes.
+Then the $Z$ lineshape scan should be used to give $\Delta\chi^{2}_{\rm min}=349$ and $696$ at CEPC and FCC-ee, respectively.
 Fermi Constant and $WW$ Threshold Scan – The Fermi constant $G_{F}$ has been measured very precisely in processes such as muon decay Workman:2022ynf .
@@
 $\chi^{2}_{W}$ $\equiv\left(\frac{m_{W,\rm th}/R_{W}-m_{W}}{\Delta m_{W}}\right)^{2}+\left(\frac{\Gamma_{W,\rm th}R_{W}-\Gamma_{W}}{\Delta\Gamma_{W}}\right)^{2}$
-$+\left[\frac{G_{F,\rm th}-G_{F}}{\Delta G_{F}}\right]^{2},$ (16)
+$+\left(\frac{G_{F,\rm th}-G_{F}}{\Delta G_{F}}\right)^{2},$ (16)
 where $R_{W}$ takes the similar form as $R_{Z}$ in Eq.
@@
 The sensitivity can reach $\Delta\chi^{2}_{W,\rm min}=169$, 263, and 7.3 at CEPC, FCC-ee, and ILC, respectively.
-By fitting the $WW$ threshold scan data, the extract $W$ mass values from the Breit-Wigner and theoretical schemes should also follow the scheme conversion in Eq.
+By fitting the $WW$ threshold scan data, the extracted $W$ mass values from the Breit-Wigner and theoretical schemes should also follow the scheme conversion in Eq.
 (10).
-However, the precisely measurement Fermi constant tends to give exactly the same value for $m_{W}$ in both schemes.
-This contradiction gives the distinguishing power.
+However, the precise measurement of the Fermi constant tends to give exactly the same value for $m_{W}$ in both schemes.
+These different scaling behaviors provide the distinguishing power.
 Conclusions – Although various resonance schemes of unstable particles have been extensively studied, using electroweak precision measurement data to verify which scheme is correct is still an open question.